Please help me with this simultaneous equation
x+2y=5
3x+y=5

Answer by tutorcecilia(2152)

x+2y=5
-2(3x+y=5) [multiply through by (-2) in order to eliminate the y term]
_______


.
.
+1x+2y=5
-6x-2y=-10
_________
-5x+0y=-5
.
-5x=-5 [solve for x]
x=1 [plug into either original equation and solve for y]
.
.
x+2y=5
(1)+2y=5
1+2y=5
2y=4
y=4/2
y=2
.
Check by plugging in the x and y values into either equation.
3x+y=5
3(1)+(2)=5
3+2=5
5=5 [checks out]